OpenStudy (anonymous):

Integration! \[\int x^2\sin x dx\]

OpenStudy (anonymous):

I am guessing this is a fairly run of the mill integration by parts problem. I'll proceed that way unless someone tells me differently.

OpenStudy (anonymous):

Yup,by parts

OpenStudy (anonymous):

\[\LARGE x^2 \int\limits \sin xdx - \int\limits( \frac{d}{dx}x^2 \int\limits \sin xdx )dx\]

OpenStudy (anonymous):

You can solve now,by parts twice fairly simple.

OpenStudy (anonymous):

I sort of see what you're doing but I think that's a bit too streamlined for my current level of (lack of) confidence with integration by parts. But anyway, here's what I've got so far. \[u = x^2 \\ du = 2x\]\[dv = \sin x \\ v = \cos x\] \[x^2\cos x - 2 \int x\cos x dx\]

OpenStudy (anonymous):

Wait. \(-\cos x\) \[-x^2\cos x + 2\int x \cos x dx\]

OpenStudy (anonymous):

yup correct,now do same for x cosx

OpenStudy (anonymous):

\[u = x \\ du = dx\]\[dv = \cos x \\ v = \sin x\] So that gives me \[-x^2\cos x + 2x\sin x - \int \sin x dx\] One more. \[-x^2\cos x + 2x\sin x + \cos x + C\] Huge success?

OpenStudy (anonymous):